OpenStudy (anonymous):

A single fair die is tossed. Find the probability of rolling a number greater than 1. * What is the probability of rolling a number greater than 1​?

OpenStudy (anonymous):

S={1,2,3,4,5,6} E={2,3,4,5,6} <greater than 1 P(E)= \[\frac{ n(E) }{ n(S) }\]

OpenStudy (superdavesuper):

Correct - u can get the ans from there @Novice3000

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }\] ?

OpenStudy (superdavesuper):

errr what is n(E)?

OpenStudy (anonymous):

I know S=6 but im not sure what E is

OpenStudy (superdavesuper):

E is when u roll a fair dice n get a number >1 n u have listed them above already...

OpenStudy (anonymous):

ok do i just choose one or use a specific number greater than 1?

OpenStudy (solomonzelman):

You have 6 possible outcomes (1-6 on the dice), and 5 "desirable" (or "wanted"), outcomes (2-6 that are greater than 1).

OpenStudy (solomonzelman):

So what is the chance/probability that out of 6 possible outcomes it will lay on one of the 5 that you want?

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\]

OpenStudy (solomonzelman):

For example, if I have 20 marbles - 15 blue marbles and 5 red marbles. What is the probability that a randomly selected marble is blue? (assuming that each marble is equally likely to be selected) ANSWER: P{blue marble} = 15/20 = 3/4

OpenStudy (anonymous):

5/6 dude

OpenStudy (solomonzelman):

Now, You have 6 possible outcomes on the dice roll. Five outcomes are good (2 through 6), and one outcome is bad (rolling "1" ). What is the probability that your outcome is good?